A hollow circular section having external dia 300mm and internal dia 250mm carries a load of 100kN at an eccentricity of 125mm. What are the maximum and minimun stess in the section.A hollow circular section having external dia 300mm and internal dia 250mm carries a load of 100kN at an eccentricity of 125m
Answer #1

stuess at A J & J 2 JA= 4.63 +9.108 13.738 N/mm² (compressive) T-02 Stuess at B JB = = 4.63-9.108 - 4.478 N/mm² (tencile)
Answer #2
13.73N/mm2 at A -4.47N/mm2 at B
source: Solid mechanics
answered by: anonymous